OpenStudy (anonymous):

what is the standard form of equation line passing through (5,-3) with a slope of -3

OpenStudy (nikita2):

Ok, y = 3x + C -- line with slop of 3. Let's find C: -3 = 3*5 + C => C = -18. And the standard form is y = 3x - 18.

OpenStudy (anonymous):

A linear equation is written in general terms y = mx + c Where m is the gradient of the line or its slope and c is the point of intercetion on the y axis. Given the coordinate pair and the slope substitute for m, x and y and start to build the equation. y = mx + c becomes... -3 = (-3)(5) + c therefore... -3 = -15 + c adding 15 to both sides... 12 = c And so the equation for this line is... y = 12 -3x

OpenStudy (anonymous):

Nikita I think you forgot the negative in the value of m.